# Conservative Closures of the Vlasov-Poisson Equations Discretized with a Symmetrically Weighted Hermite Spectral Expansion in Velocity

### v1

@@ Acknowledgement
O.I. and B.K. were partially supported by the National Science Foundation under Award 2028125 for “SWQU: Composable Next Generation Software Framework for Space Weather Data Assimilation and Uncertainty Quantification”. O.K., G.L.D., and O.I. were supported by the Laboratory Directed Research and Development Program of Los Alamos National Laboratory under project number 20220104DR. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). F.D.H. was supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, Theory Program, under Award No. DE-FG02-95ER54309.

### v2

@@ Acknowledgement
O.I. was partially supported by the Los Alamos National Laboratory (LANL) Student Fellowship sponsored by the Center for Space and Earth Science (CSES). CSES is funded by LANL’s Laboratory Directed Research and Development (LDRD) program under project number 20210528CR. O.I. and B.K. were partially supported by the National Science Foundation under Award 2028125 for “SWQU: Composable Next Generation Software Framework for Space Weather Data Assimilation and Uncertainty Quantification”. O.K., G.L.D., and O.I. were supported by the LDRD Program of LANL under project number 20220104DR. LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). F.D.H. was supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, Theory Program, under Award No. DE-FG02-95ER54309.